# Predicted High-Pressure Hot Superconductivity in Li2CaH16 and Li2CaH17 Phases that Resemble the Type-II Clathrate Structure

### v1

@@ Abstract
High-temperature high-pressure superconducting hydrides are typically characterized by cage-like hydrogenic lattices filled with electropositive metal atoms. Here, density functional theory based evolutionary crystal structure searches find two phases that possess these geometric features and are related to the Type-II clathrate structure. In these Li2CaH16 and Li2CaH17 phases the calcium atom occupies the larger cage and the lithium atom the smaller one. The highest superconducting critical temperatures predicted within the isotropic Eliashberg formalism, 330 K at 350 GPa for Li2CaH16 and 370 K at 300 GPa for Li2CaH17, suggest these structures fall in the class of high-energy-density quantum materials known as hot superconductors. As pressure is lowered the cage-like lattices distort with the emergence of quasimolecular hydrogenic motifs; nonetheless Li2CaH17 is predicted to be superconducting down to 160 GPa at 205 K.
@@ 1 Introduction
In 1968, Neil Ashcroft predicted that metallic hydrogen – the hypothetical ground state of the lightest element under pressure – could behave as a high-temperature superconductor 1. Nearly four decades later Ashcroft described how the pressures required to metallize hydrogen could be decreased via addition of a second element that "chemically precompressed" the system 2. As early as 2015, experimental studies supporting Ashcroft’s prediction came to the fore. These included measurements of high superconducting critical temperatures (Tcs) in phases with unique stoichiometries, stable only under pressure, including H3S (Tc = 203 K, 150 GPa) 3, LaH10 (Tc = 260 K, 200 GPa) 4, 5, YH9 (Tc = 262 K, 182 GPa) 6, YH6 (Tc = 224 K, 166 GPa) 7, CaH6 (Tc = 210-215 K, 160-172 GPa) 8, 9, 10, and CeH9 11, 12, 13 (Tc = 100 K, 166 GPa). Recently, local magnetometry experiments on CeH9, performed via nitrogen-vacancy centers embedded within the diamond anvil, provided evidence of the Meissner effect 14.
@@
The immense pressures required to stabilize the superconducting binary hydrides can only be reached in diamond anvil cells and come coupled with experimental intricacy and expense, making theoretical calculations an appealing method to both determine systems of interest and interpret experimental results 15, 16. Density functional theory (DFT) calculations foster accurate predictions for superconducting properties (if properly converged), and crystal structure prediction (CSP) techniques have been successful in uncovering promising synthetic targets (for unit cells that are not too complex) 17. The chemical space has been thoroughly scoured for fruitful high-temperature superconducting binary hydrides 18, 19, 20. In contrast, ternary hydrides present relatively unexplored possibilities, however predictions for this class of hydrides are significantly more challenging 21. The vast chemical space means that an increased number of stoichiometries must be considered for each elemental combination, with a more complex potential energy landscape that requires a greater number of structures be optimized in the CSP runs 22. To overcome these challenges, machine learning potentials have been used to roughly explore the phase diagram prior to performing more costly DFT optimizations 23, and high-throughput calculations have been employed to study elemental substitutions on previously-discovered superconducting templates 24. The introduction of a third element could potentially be used to lower the stabilization pressure, such as in synthesized LaBeH8 (Tc = 110 K, 80 GPa) 25 and predicted BaSiH8 (Tc = 71 K, 3 GPa) 24, or increase the superconducting critical temperature as in the synthesized (La,Ce)H9 (Tc = 148-178 K, 97-172 GPa) 26 and predicted Li2ScH20 (Tc = 242 K, 300 GPa) 27. The former strategy is usually accomplished by using one electropositive element and one -block element, whereas the latter strategy usually requires the combination of two electropositive elements. Recently, superconductivity in an entropy-stabilized quaternary hydride, (La,Y,Ce)H10, has been reported with a maximum Tc of 190 K at 112 GPa 28.
@@
The hydrogenic lattices of the binary hydride superconductors of an electropositive element that have very high s generally assume cage-like structures within which a metal atom is encapsulated. Several stoichiometries are well known. The MH6 sodalite-like clathrate, first exemplified by CaH6 8, 9, 10, has a metal atom, M, centered in an H24 cage consisting of six square and eight hexagonal faces. The MH9 structure, which surrounds the metal atom with an H29 cage of six irregular squares, six pentagons and six hexagons, is found in numerous rare-earth compounds including synthesized YH9 6 and CeH9 11, 12, 14. Further increasing the hydrogen content, MH10 stoichiometries, where the metal atom sits within hydrogenic polyhedra of six squares and twelve hexagons, are perhaps the most well-known owing to the synthesis of the LaH10 4, 5 superhydride. The structure of the hydrogen lattice in these MH10 systems is known as “AST” in the clathrate and zeolite community 29. More recently, larger hydride cages have been predicted to encapsulate rare earth metals. These include MH18 (M=Y, La, Ce, Ac, Th) stoichiometries with cubic ( and ) space groups, possessing H36 cages, with a maximum predicted Tc of 329 K at 350 GPa for CeH18 30. Can we draw inspiration from these systems and other occurring clathrates, specifically with phases that contain two electropositive elements?
@@
Looking across the periodic table to other cage-forming elements we encounter silicon, which is able to form large polyhedral structures of sp3 bonded Si atoms that are filled with metal atoms, characterizing the silicon clathrates (SiCL). Two of the most well studied SiCLs are the type-I M8Si24 (the Weaire-Phelan structure), which can be interpreted as two Si20 and 6 Si24 cages filled with a metal atom, and the type-II Mx≤24Si136 structure, which is made of 16 Si20 and 8 Si28 cages with a variable number of metal atoms that can be controlled by the synthesis conditions 31, 32, 33, 34, 35. Several of the known silicon clathrates are superconductors, for example the type-I Ba8Si46 36, with an ambient pressure Tc of 8 K. Recent experiments report the synthesis of (Ba/La/Eu)4H23 37, 38, 39 high pressure hydrides in the type-I structure, with untested Tcs, and Lu4H23 40, with a Tc of 73 K at 218 GPa. Furthermore, predicted Li2(La/Y)H17 27 both match the type-II structure, with the highest Tc being 156 K at 160 GPa for Li2LaH17. These structure types have also been explored for Li2NaH16, Li2NaH17, and LiNa3H23 41, predicting a maximum Tc of 340 K at 300 GPa for Li2NaH17. Prior to these structural correlations being drawn, a computational study predicted that when the MgH16 42 structure was doped with Li, an Li2MgH16 “hot superconductor” 43, which differs from the type-II structure by only a single Wyckoff site, possessing a record calculated maximum Tc of 473 K at 250 GPa could be stabilized 43.
@@
Herein, through careful DFT calculations we investigated if a high-temperature superconductor analogous to Li2MgH16 43, but where Mg is replaced by Ca, could be stabilized under pressure. Because the type-I and type-II SiCL structures each possess two cages with distinct sizes, we hypothesized they would be particularly well suited templates for incorporating mixed metal atoms, and in our specific case lithium would fill the smaller cages and calcium the larger ones. Lithium is light, lending itself to strong electron-phonon coupling, with the goal of increasing Tc. Furthermore, as several Ca hydrides are synthesized or predicted superconductors, we intended to improve upon their properties with the addition of Li. With these motivations, we performed CSP searches at megabar pressures for Li2CaHn (). Both stable and metastable phases were discovered, and their electronic structures were analyzed. Two dynamically stable metallic structures, one that is a distortion of the type-II SiCL, and one that can be derived from it, were found, and electron-phonon coupling calculations were performed to estimate their isotropic Eliashberg Tcs: 330 K at 350 GPa for Li2CaH16, and 205 K at 160 GPa or 370 K at 300 GPa for Li2CaH17. This work extends the list of predicted “hot” superconducting ternary hydrides by two.
@@ 2 Methods
Exploratory crystal structure prediction (CSP) searches using the open-source XTALOPT 44, 45 evolutionary algorithm (EA) 12th 46 release were performed to identify stable and metastable phases. EA runs were performed for the following stoichiometries: Li2CaHn (=10-20) and Li2Ca at 100, 200, and 300 GPa. The initial generation consisted of random symmetric structures created by the RANDSPG algorithm.47 Duplicate structures were identified via the XTALCOMP algorithm for removal from the breeding pool.48 The breeding operators employed and their relative probabilities were: (i) stripple (50%), (ii) permustrain (35%), and (iii) crossover (15%). The minimum number of generated structures was set to 1500. Each structure optimization consisted of four consecutive geometry optimizations, with each increasing in precision. The final geometry optimization step of the EA search employed a cutoff energy of 400 eV and a -grid generated using the -centered Monkhorst-Pack scheme 49, with the number of divisions along each reciprocal lattice vector chosen such that the product of this number with the real lattice constant was 30 Å. After the EA searches concluded, the three most enthalpically favored unique structures at each stoichiometry and pressure were used for further analysis.
@@
Precise geometry optimizations, electronic structure calculations, electron localization functions (ELF), Bader charges, and first principles molecular dynamics (FPMD) simulations were performed using DFT calculations with the PBE functional as implemented in the Vienna Ab Initio Simulation Package (VASP). 50, 51 We employed a planewave basis set with an energy cutoff of 600 eV, along with PAW potentials 52 where the H 1s1, Li 1s22s22p0 and Ca 3p64s23d0.01 electrons were treated explicitly in all calculations. -meshes were generated using the -centered Monkhorst-Pack scheme 49, and the number of divisions along each reciprocal lattice vector was chosen such that the product of this number with the real lattice constant was 50 Å. All FPMD simulations were performed with a 2 2 2 supercell, a single -point at , an ensemble (constant number of particles, constant volume, constant temperature), Nosé-Hoover thermostat 53, a 0.5 femtosecond timestep, and temperatures corresponding to the predicted s. The crystal orbital Hamilton populations (COHPs) and the negative of the COHP integrated to the Fermi level (-iCOHP) were calculated using the LOBSTER package to analyze bonding. 54, 55
@@
To determine if the predicted structures were dynamically stable, and obtain their zero-point energy, harmonic phonons were calculated using the supercell approach. In the supercell approach, Hellmann-Feynman forces were calculated from a supercell constructed by replicating the optimized structure wherein the atoms had been displaced, and dynamical matrices were computed using the PHONOPY code. 56 Quantum ESPRESSO (QE) was used to obtain the dynamical matrix and electron-phonon coupling (EPC) parameters.57 H, Li, and Ca pseudopotentials obtained from the QE pseudopotential library were generated by the Vanderbilt ultrasoft method with a 1s1 configuration for H, a 1s22s12p0 configuration for Li, and a 3s23p64s23d04p0 configuration for Ca. The PBE functional was employed for all QE calculations. Plane-wave basis set cutoff energies were set to 70 Rydbergs (Ry) for all systems. The Methfessel-Paxton Brillouin-zone smearing of 0.015 Ry was applied.58 The EPC parameter was calculated using Gaussian broadenings in increments of 0.005 Ry ranging from 0.005 to 0.050 Ry. A 15 15 15 -grid with a -mesh of 5 5 5 was employed for Li2CaH17 at 300 GPa. Subsequent calculations on Li2CaH17 utilized a -mesh of 3 3 3, which was found to be sufficiently converged for EPC properties (Table S7). Therefore, a 15 15 15 -grid with a q-mesh of 3 3 3 was utilized for Li2CaH16. For Li2CaH17 the EPC appeared to be converged with these parameters at a broadening of 0.030 Ry, and Li2CaH16 converged at a broadening of 0.050 Ry. Critical temperature (Tc) values were obtained by numerical solutions of the Eliashberg equation 59 with a Coulomb repulsion parameter, , of 0.1.
@@ 3.1 Hydrogenic Motifs
Let us take a closer look at the predicted Li2CaHn structures, to better understand the peculiarities of their atomic configurations, and in particular the hydrogenic motifs that are so important for their superconducting properties. Based on the extensive work performed on binary high-pressure hydrides containing electropositive elements 20, 22, it is known that the presence of H- and H anions is detrimental for superconductivity. When these are present, the hydrides are metallized via band induced broadening and overlap, resulting in a low density of states (DOS) at the Fermi level (). Typically, compounds with H motifs are predicted to possess intermediate s owing to the metallicity that stems from the partial filling of the antibonding H2 orbitals, and a comparatively high DOS at (). However, increasing the dimensionality of the hydrogenic lattices, from 1D chains to 2D sheets to 3D extended lattices, maintains a high , and increases the number of phonon modes that can be involved in the EPC mechanism, thereby resulting in the highest predicted s. For this reason, compounds with cage-like lattices that enclathrate electropositive atoms are among those with the highest computed and measured superconducting critical temperatures.
@@
Analysis of the structures found in our CSP searches reveals these same hydrogenic motifs, and illustrates how they vary as a function of hydrogen content and pressure (Table S8). The Li2CaHn stoichiometries were typically characterized by H- and bent, asymmetric H units, interspersed with H2 molecules at all pressures investigated. The most stable phases featured these same motifs at 100 and 200 GPa, however by 300 GPa incomplete clathrate cages began to emerge, suggesting that both the metal to hydrogen ratio and pressure are key for clathrate formation in the high-pressure hydrides. A clathrate cage is found within Li2CaH16 at 300 GPa, but it dissociates to an incomplete clathrate at 200 GPa and further to a hydrogenic lattice containing discrete H, H2 and H- units by 100 GPa. The global minimum Li2CaH17 phase possesses a clathrate structure at both 300 and 200 GPa, but dissociates to form H2 and H- by 100 GPa. Within Li2CaH18 and Li2CaH19 incomplete clathrates are present at 300 GPa, with H2 units interspersed within the lattice for . At pressures of 100 and 200 GPa, these incomplete clathrates dissociate to form H, H2 and H-. Lastly, Li2CaH20 possesses only H2 units at 100, 200, and 300 GPa. Our observations are in-line with a chemical pressure analysis, which illustrates that clathrate formation in the high-pressure hydrides depends on the size-match between the metal cation and the hydrogenic cluster encapsulating it 60. We postulated that the Li2CaH16 and the Li2CaH17 structures could be candidates for high-temperature, and perhaps “hot” superconductivity, given the proclivity of clathrate-structure-types for large EPC values.

### v2

@@ Abstract
High-temperature high-pressure superconducting hydrides are typically characterized by cage-like hydrogenic lattices filled with electropositive metal atoms. Here, density functional theory based evolutionary crystal structure searches find two phases that possess these geometric features and are related to the Type-II clathrate structure. In these Li2CaH16 and Li2CaH17 phases the calcium atom occupies the larger cage and the lithium atom the smaller one. The highest superconducting critical temperatures predicted within the isotropic Eliashberg formalism, 330 K at 350 GPa for Li2CaH16 and 370 K at 300 GPa for Li2CaH17, suggest these structures are another example of hot superconducting hydrides. As pressure is lowered the cage-like lattices distort with the emergence of quasimolecular hydrogenic motifs; nonetheless Li2CaH17 is predicted to be superconducting down to 160 GPa at 205 K.
@@ 1 Introduction
In 1968, Neil Ashcroft predicted that metallic hydrogen – the hypothetical ground state of the lightest element under pressure – could behave as a high-temperature superconductor 1. The realization of this state has proven elusive, with the most recently estimated pressure needed being over 570 GPa 2. Nearly four decades later Ashcroft described how the pressures required to metallize hydrogen could be decreased via addition of a second element that "chemically precompressed" the system 3. As early as 2015, experimental studies supporting Ashcroft’s prediction came to the fore. These included measurements of high superconducting critical temperatures (s) in phases with unique stoichiometries, stable only under pressure, including H3S ( = 203 K, 150 GPa) 4, LaH10 ( = 260 K, 200 GPa) 5, 6, YH9 ( = 262 K, 182 GPa) 7, YH6 ( = 224 K, 166 GPa) 8, CaH6 ( = 210-215 K, 160-172 GPa) 9, 10, 11, and CeH9 12, 13, 14 ( = 100 K, 166 GPa). Recently, local magnetometry experiments on CeH9, performed via nitrogen-vacancy centers embedded within the diamond anvil, provided evidence of the Meissner effect 15.
@@
The immense pressures required to stabilize the superconducting binary hydrides can only be reached in diamond anvil cells and come coupled with experimental intricacy and expense, making theoretical calculations an appealing method to both determine systems of interest and interpret experimental results 16, 17. Density functional theory (DFT) calculations have been found capable of predicting superconducting properties of conventional superconductors, and crystal structure prediction (CSP) techniques have been successful in uncovering promising synthetic targets (for unit cells that are not too complex) 18. The chemical space has been thoroughly scoured for fruitful high-temperature superconducting binary hydrides 19, 20, 21. In contrast, ternary hydrides present relatively unexplored possibilities, however predictions for this class of hydrides are significantly more challenging 22. The vast chemical space means that an increased number of stoichiometries must be considered for each elemental combination, with a more complex potential energy landscape that requires a greater number of structures be optimized in the CSP runs 23. To overcome these challenges, machine learning potentials have been used to roughly explore the phase diagram prior to performing more costly DFT optimizations 24, and high-throughput calculations have been employed to study elemental substitutions on previously-discovered superconducting templates 25. The introduction of a third element could potentially be used to lower the stabilization pressure, such as in synthesized LaBeH8 ( = 110 K, 80 GPa) 26 and predicted BaSiH8 ( = 71 K, 3 GPa) 25, or increase the superconducting critical temperature as in the synthesized (La,Ce)H9 ( = 148-178 K, 97-172 GPa) 27 and predicted Li2ScH20 ( = 242 K, 300 GPa) 28. The former strategy is usually accomplished by using one electropositive element and one -block element, whereas the latter strategy usually requires the combination of two electropositive elements. Recently, superconductivity in a quaternary hydride, (La,Y,Ce)H10, has been reported with a maximum of 190 K at 112 GPa 29.
@@
The hydrogenic lattices of the binary hydride superconductors of an electropositive element that have very high s generally assume cage-like structures within which a metal atom is encapsulated. Several stoichiometries are well known. The MH6 sodalite-like clathrate, first exemplified by CaH6 9, 10, 11, has a metal atom, M, centered in an H24 cage consisting of six square and eight hexagonal faces. The MH9 structure, which surrounds the metal atom with an H29 cage of six irregular squares, six pentagons and six hexagons, is found in numerous rare-earth compounds including synthesized YH9 7 and CeH9 12, 13, 15. Further increasing the hydrogen content, MH10 stoichiometries, where the metal atom sits within hydrogenic polyhedra of six squares and twelve hexagons, are perhaps the most well-known owing to the synthesis of the LaH10 5, 6 superhydride. The structure of the hydrogen lattice in these MH10 systems is known as “AST” in the clathrate and zeolite community 30. More recently, larger hydride cages have been predicted to encapsulate rare earth metals. These include MH18 (M=Y, La, Ce, Ac, Th) stoichiometries with cubic ( and ) space groups, possessing H36 cages, with a maximum predicted of 329 K at 350 GPa for CeH18 31. Can we draw inspiration from these systems and other occurring clathrates, specifically with phases that contain two electropositive elements?
@@
Looking across the periodic table to other cage-forming elements we encounter silicon, which is able to form large polyhedral structures of sp3 bonded Si atoms that are filled with metal atoms, characterizing the silicon clathrates (SiCL). Two of the most well studied SiCLs are the type-I M8Si24 (the Weaire-Phelan structure), which can be interpreted as two Si20 and 6 Si24 cages filled with a metal atom, and the type-II Mx≤24Si136 structure, which is made of 16 Si20 and 8 Si28 cages with a variable number of metal atoms that can be controlled by the synthesis conditions 32, 33, 34, 35, 36. Several of the known silicon clathrates are superconductors, for example the type-I Ba8Si46 37, with an ambient pressure of 8 K. Recent experiments report the synthesis of (Ba/La/Eu)4H23 38, 39, 40 high pressure hydrides in the type-I structure, with untested s, and Lu4H23 41, with a of 73 K at 218 GPa. Furthermore, predicted Li2(La/Y)H17 28 both match the type-II structure, with the highest being 156 K at 160 GPa for Li2LaH17. These structure types have also been explored for Li2NaH16, Li2NaH17, and LiNa3H23 42, predicting a maximum of 340 K at 300 GPa for Li2NaH17. Prior to these structural correlations being drawn, a computational study predicted that when the MgH16 43 structure was doped with Li, an Li2MgH16 “hot superconductor” 44, which differs from the type-II clathrate structure by a single unoccupied Wyckoff site (a clathrate atom at the (8a) 1/8, 1/8, 1/8 site), possessing a record calculated maximum of 473 K at 250 GPa could be stabilized 44.
@@
Herein, through careful DFT calculations we investigated if a high-temperature superconductor analogous to Li2MgH16 44, but where Mg is replaced by Ca, could be stabilized under pressure. Because the type-I and type-II SiCL structures each possess two cages with distinct sizes, we hypothesized they would be particularly well suited templates for incorporating mixed metal atoms, and in our specific case lithium would fill the smaller cages and calcium the larger ones. Lithium is both electropositive, donating charge to the surrounding hydrogen lattice, and light, allowing for engaging in coupled high-frequency phonons with the hydrogen lattice. Furthermore, as several Ca hydrides are synthesized or predicted superconductors, we intended to improve upon their properties with the addition of Li. With these motivations, we performed CSP searches at megabar pressures for Li2CaHn (). Both stable and metastable phases were discovered, and their electronic structures were analyzed. Two dynamically stable metallic structures, one that is a distortion of the type-II SiCL, and one that can be derived from it, were found, and electron-phonon coupling calculations were performed to estimate their isotropic Eliashberg s: 330 K at 350 GPa for Li2CaH16, and 205 K at 160 GPa or 370 K at 300 GPa for Li2CaH17. This work extends the list of predicted “hot” superconducting ternary hydrides by two.
@@ 2 Methods
Exploratory crystal structure prediction (CSP) searches using the open-source XTALOPT 45, 46 evolutionary algorithm (EA) 12th 47 release were performed to identify stable and metastable phases. EA runs were performed for the following stoichiometries: Li2CaHn (=10-20) and Li2Ca at 100, 200, and 300 GPa. The initial generation consisted of random symmetric structures created by the RANDSPG algorithm.48 Duplicate structures were identified via the XTALCOMP algorithm for removal from the breeding pool.49 The breeding operators employed and their relative probabilities were: (i) stripple (50%), (ii) permustrain (35%), and (iii) crossover (15%). The minimum number of generated structures was set to 1500. Each structure optimization consisted of four consecutive geometry optimizations, with each increasing in precision. The final geometry optimization step of the EA search employed a cutoff energy of 400 eV and a -grid generated using the -centered Monkhorst-Pack scheme 50, with the number of divisions along each reciprocal lattice vector chosen such that the product of this number with the real lattice constant was 30 Å. After the EA searches concluded, the three most enthalpically favored unique structures at each stoichiometry and pressure were used for further analysis.
@@
Precise geometry optimizations, electronic structure calculations, electron localization functions (ELF), Bader charges, and first principles molecular dynamics (FPMD) simulations were performed using DFT calculations with the PBE functional as implemented in the Vienna Ab Initio Simulation Package (VASP). 51, 52 We employed a planewave basis set with an energy cutoff of 600 eV, along with PAW potentials 53 where the H 1s1, Li 1s22s12p0 and Ca 3p64s23d0.01 electrons were treated explicitly in all calculations. -meshes were generated using the -centered Monkhorst-Pack scheme 50, and the number of divisions along each reciprocal lattice vector was chosen such that the product of this number with the real lattice constant was 50 Å. All FPMD simulations were performed with a 2 2 2 supercell, a single -point at , an ensemble (constant number of particles, constant volume, constant temperature), Nosé-Hoover thermostat 54, a 0.5 femtosecond timestep, and temperatures corresponding to the predicted s. The crystal orbital Hamilton populations (COHPs) and the negative of the COHP integrated to the Fermi level (-iCOHP) were calculated using the LOBSTER package to analyze bonding. 55, 56
@@
To determine if the predicted structures were dynamically stable, and obtain their zero-point energy, harmonic phonons were calculated using the supercell approach. In the supercell approach, Hellmann-Feynman forces were calculated from a supercell constructed by replicating the optimized structure wherein the atoms had been displaced, and dynamical matrices were computed using the PHONOPY code. 57 Quantum ESPRESSO (QE) was used to obtain the dynamical matrix and electron-phonon coupling (EPC) parameters.58 H, Li, and Ca pseudopotentials obtained from the QE pseudopotential library were generated by the Vanderbilt ultrasoft method with a 1s1 configuration for H, a 1s22s12p0 configuration for Li, and a 3s23p64s23d04p0 configuration for Ca. The PBE functional was employed for all QE calculations. Plane-wave basis set cutoff energies were set to 70 Rydbergs (Ry) for all systems. The Methfessel-Paxton Brillouin-zone smearing of 0.015 Ry was applied.59 The EPC parameter was calculated using Gaussian broadenings in increments of 0.005 Ry ranging from 0.005 to 0.050 Ry. A 15 15 15 -grid with a -mesh of 5 5 5 was employed for Li2CaH17 at 300 GPa. Subsequent calculations on Li2CaH17 utilized a -mesh of 3 3 3. This sparser -mesh resulted in a deviation in of <2%, with a 20 K reduction in , and an overall reduction in computational cost by 80%. This was determined to be sufficiently converged for prediction of EPC properties (Table S7). Therefore, a 15 15 15 -grid with a q-mesh of 3 3 3 was utilized for Li2CaH16. For Li2CaH17 the EPC appeared to be converged with these parameters at a broadening of 0.030 Ry, and Li2CaH16 converged at a broadening of 0.050 Ry.
Critical temperature () values were obtained with a Coulomb repulsion parameter, , of 0.1 by the Allen-Dynes modified McMillan equation for strong coupling,
where is the logarithmic average frequency and the and correction factors for strong coupling and shape dependence are functions of , , and (equaling unity for the weak coupling limit, see Ref. 60). s were also obtained by solving the isotropic Eliashberg equations 61 within the Fermi-surface-restricted (FSR) approximation numerically based on the spectral function, , obtained from the QE calculations. Isotropic solutions of values in hydrides have been found to underestimate s by 10% compared to anisotropic solutions 62, 63, 64, 65, and a recent implementation has made it possible to compare the results of the full bandwidth (FBW) approach to those obtained with the FSR approach for hydrides66. The Matsubara frequency cutoff was set to a value of 20 , the root-mean-square phonon frequency, while the number of frequencies was allowed to vary up to 40,000.
@@ 3.1 Hydrogenic Motifs
Let us take a closer look at the predicted Li2CaHn structures, to better understand the peculiarities of their atomic configurations, and in particular the hydrogenic motifs that are so important for their superconducting properties. Based on the extensive work performed on binary high-pressure hydrides containing electropositive elements 21, 23, it is known that the presence of H- and H anions is detrimental for superconductivity. When these molecules are present, the hydrides are metallized via pressure-induced broadening of the H- and H bonding and anti-bonding bands, resulting in a low density of states (DOS) at the Fermi level () and no or low 67, 68. Typically, compounds with H motifs are predicted to possess intermediate s owing to the metallicity that stems from the partial filling of the antibonding H2 orbitals, and a comparatively high DOS at (). Expanding from molecules into extended hydrogenic lattices, a high can be maintained. Compounds with cage-like lattices that enclathrate electropositive atoms are found to possess large values of , the electron-phonon coupling constant, leading to the highest computed and measured superconducting critical temperatures.69
@@
Analysis of the structures found in our CSP searches reveals these same hydrogenic motifs, and illustrates how they vary as a function of hydrogen content and pressure (Table 1). The Li2CaHn stoichiometries were typically characterized by H- and bent, asymmetric H units, interspersed with H2 molecules at all pressures investigated. The most stable phases featured these same motifs at 100 and 200 GPa, however by 300 GPa incomplete clathrate cages began to emerge, suggesting that both the metal to hydrogen ratio and pressure are key for clathrate formation in the high-pressure hydrides. A clathrate cage is found within Li2CaH16 at 300 GPa, but it dissociates to an incomplete clathrate at 200 GPa and further to a hydrogenic lattice containing discrete H, H2 and H- units by 100 GPa. The global minimum Li2CaH17 phase possesses a clathrate structure at both 300 and 200 GPa, but dissociates to form H2 and H- by 100 GPa. Within Li2CaH18 and Li2CaH19 incomplete clathrates are present at 300 GPa, with H2 units interspersed within the lattice for . At pressures of 100 and 200 GPa, these incomplete clathrates dissociate to form H, H2 and H-. Lastly, Li2CaH20 possesses only H2 units at 100, 200, and 300 GPa. Our observations are in-line with a chemical pressure analysis, which illustrates that clathrate formation in the high-pressure hydrides depends on the size-match between the metal cation and the hydrogenic cluster encapsulating it 70. We postulated that the Li2CaH16 and the Li2CaH17 structures could be candidates for high-temperature, and perhaps “hot” superconductivity, given the proclivity of clathrate-structure-types for large EPC values.
Stoichiometry | 100 GPa | 200 GPa | 300 GPa |
Li2CaH10 | H,H2,H- | H2,H- | H,H2,H- |
Li2CaH11 | H2,H- | H2,H- | H2 |
Li2CaH12 | H2,H- | H2,H- | H,H2,H- |
Li2CaH13 | H2,H- | H2,H- | H2 |
Li2CaH14 | H2,H- | H2,H- | Incomplete Clathrate |
Li2CaH15 | H,H2,H- | H2,H- | Incomplete Clathrate |
Li2CaH16 | H,H2,H- | Incomplete Clathrate | Clathrate |
Li2CaH17 | H2,H- | Clathrate | Clathrate |
Li2CaH18 | H2,H- | H2,H- | Incomplete Clathrate, H2 |
Li2CaH19 | H,H2,H- | H,H2,H- | Incomplete Clathrate |
Li2CaH20 | H2 | H2 | H2 |